OpenStudy (anonymous):

The width of a rectangle is 7 cm shorter than its length and its area is 228 sq. cm. Find the dimensions of the rectangle

OpenStudy (angela210793):

A=l*w=l*(l-7)=228 l^2-7l-228=0 now solve it by using quadratic formula